# B-VLLM: A Vision Large Language Model with Balanced Spatio-Temporal Tokens

### v1

@@ 5 Experimental Result & Discussion
$\Delta$ +2.7 +1.0 +0.6 +46 0.0 -0.2 +2.8 +1.8 +2.8
B-VLLM 32 71.8 83.7 46.5 1504 76.6 74.5 59.7 54.2 67.1
@@ 5.1 Video Benchmark Quantitative Evaluation
We evaluate the proposed B-VLLM on various video benchmarks in a zero-shot manner. The video duration in the these benchmarks spans from 10 second to 1 hour, where the details are summarized in the supplementary materials. For open-end answering, we have MSVD-QA [43], MSRVTT-QA [43] and ActivityNet-QA [9], which aim to evaluate the capacity of video captioning. Following recent works [32, 6, 24], for benchmarking with open-end question answering, performance is evaluated by comparing the similarity between the outputs of VLLMs and ground truth text using proprietary LLMs.
@@
For multi-choice question answering (MCQA), we have MVBench [21] and VideoMME [12], which are multi-purpose video understanding benchmarks; EgoSchema [33] focuses on evaluating the temporal understanding capability; Perception [34] aims to examine reasoning and perception skills of VLLMs; and VNBench [47] evaluates a VLLM using the needle in a haystack manner: asking questions about one or more irrelevant images that are randomly inserted into a video. For MCQA, the performance is evaluated based on the accuracy of selecting the correct options.
@@
Quantitative results for open-end question answering and MCQA benchmarks are reported in Table 1. Overall, B-VLLM achieves SOTA performance on MSVD-QA, Perception, VNBench, Egoschema, and VideoMME. Specifically, when integrating with existing VLLMs, it significantly boosts performance to the two baselines across most video benchmarks. Notably, our method achieves substantial improvements in long-video scenarios. For example, VideoMME-Long evaluates VLLMs performance in extremely long-video settings, with an average video duration of 2466.7 seconds. Our proposed method introduced an 10.5% performance gain for LLaMA-VID [24] and 3.8% for VideoLLaMA2 [6]. Although LLaMA-VID encodes each frame into only two visual tokens, it struggles with extremely long video due to its 2,048 context window limit. Our adaptive frame selection module addresses this limitation by selecting a fixed number of frames most relevant to the given context, ensuring the number of visual tokens remains within the context window of LLaMA-VID. As a result, LLaMA-VID consistently performs well in medium and long video settings. We also observe a noticeable performance drop when VideoLLaMA2 handles medium and long videos, which we attribute to its uniform frame sampling strategy that may overlook the frames relevant to the context. The integration of our method mitigates this issue by filtering out irrelevant frames, leading to 1.7% and 3.8% performance gains for VideoLLaMA2.
@@
Additionally, the proposed method remains effective in the short-video settings. For example, LLaMA-VID achieves a 4.5% and 5.8% performance gain in MVBench and Perception through the integration of our method. Similarly, in open-end benchmarks, performance improvements of 1.7%, 0.2% and 1.5% are observed in MSVD-QA, MSRVTT-QA and ActivityNet-QA, respectively. Such significant improvement owes to the adaptivity in frame-wise visual tokens, as the loss of spatial cues in the original LLaMA-VID may limit the VLLM’s reasoning and perception capabilities. Similar performance improvements are also observed in Video-LLaMA2, with gains of 1.0%, 2.4%, 1.8%, 2.1%, 1.9%, 3.3% and 0.9% on MVBench, Perception, VNBench, Egoschema, MSVD-QA, MSRVTT-QA and ActivtiyNet-QA, respectively.
@@
An interesting observation is that VLLMs using uniform sampling strategy tend to achieve better performance in short-video settings, especially when the number of frames sampled closely matches the video duration at 1 fps. For example, VideoChat2 [21] achieves the best accuracy in MVBench; and LLaVA-Next-Video [46] achieves the second best performance on Perception [34]. These two models sample 16 and 32 frames from each video, respectively, with the average duration of MVBench and Perception being 16 frames and 23 frames at 1 fps. These methods could achieve better performance in long-video setting if more frames are sampled. However, due to context window limitations, the number of visual tokens would quickly become overwhelming, leading to degraded performance and a heavy computation burden. In contrast, our proposed method performs consistently in both short- and long-video settings, benefiting from the frame selection and spatio-temporal selection and merging techniques.
@@ 5.2 Image Benchmark Quantitative Evaluation
Our proposed method yields comparable, if not superior, performance, despite not being specifically designed for the image-only setting and using significantly fewer visual tokens per image than existing image-based VLLMs. In particular, B-VLLM outperforms classic image-only VLLM, LLaVA-1.5 [27] on MMBench [28], ScienceQA-Image [30], and SEED-Image [19] by 7.1%, 7.7%, and 1.0%, respectively. Notably, comparable results are observed on POPE [23] and MME [10], despite LLaVA-1.5 [27] consuming 18 times more visual token than ours. These benchmarks primarily focus on reasoning and general image understanding, where fine-grained visual details are less critical and typically require fewer visual tokens. However, as noted by LLaMA-VID [24], more visual tokens generally lead to better performance on certain image-only benchmarks. The limitation in the number of visual tokens compromises our method’s performance on VizWiz [14], VQAv2 [13] and TextVQA [38], where these benchmarks examine spatial reasoning and OCR. For example, TextVQA [38] focuses on detecting text within image, while GQA [16] emphasizes complex scene understanding.
In terms of performance gain, integrating our proposed method results in significant improvements for LLaMA-VID [24] and VideoLLaMA2 [6] on image benchmarks. In particular, for LLaMA-VID, we observe gains of 9.5% and 10.2% on MMB and VQAV2; For VideoLLaMA2, non-trivial gains can be observed when incorporating with our method. We believe these gains stem from the effectiveness of the visual token sampling module, as it dynamically selects the most informative tokens, reducing the overall token count while maintaining performance.
Table 3: Ablation study on adaptive frame selection (F.S), temporal frame token merging (T.M.), and spatial visual token sampling (S.S.).
 F.S. T.M. S.S. MVB VMME MMB POPE
@@
(a) MVBench
(b) VideoMME
(c) Image Benchmarks
Figure 3: Impacts on the number of frames selected and spatial visual tokens.
@@ 5.3 Ablation Study
The ablation study is divided into two parts. First, we study the role of each module in the proposed method. Second, we assess the robustness of our method against variations in the number of frame selection and spatial visual tokens.
@@
Effectiveness of the proposed modules in terms of their performance is reported in Table 3. In this experiment, LLaMA-VID [24] is selected as the base model. MVBench [21] and VideoMME [12] are used to evaluate video understanding performance, while MMBench [28] and POPE [23] are used to assess image understanding performance. Starting with the baseline, without the integration of our proposed modules, LLaMA-VID [24] displays moderate performance on the short-video benchmark MVBench [21]. However, compromised performance can be observed on the medium-to-long video benchmark Video-MME [12] due to the limitation of the context window length. In addition, LLaMA-VID [24] typically performs worst on image benchmarks, due to its poor spatial perception caused by a deficit of spatial visual tokens.
With the introduction of the text-conditioned adaptive frame selection (F.S.) module, the baseline model no longer suffers from its own token selection mechanism, resulting in 2.9% performance boost on the medium-to-long video benchmark VideoMME [12]. With the temporal frame token merging (T.M.) mechanism, the model mitigates temporal redundancy in duplicated frame selection, resulting in noticeable performance gains on both short and medium-to-long video benchmarks. Similarly, the integration of spatial visual token sampling (S.S.) enables dynamic adjustment on the number of spatial visual tokens to balance the temporal and spatial cues, resulting in performance gains 1.4% and 1.5% on MVBench and VideoMME, respectively.
@@
Impacts on the number of frame selected and spatial visual tokens are illustrated in Figure 3. In our experiments, we examine the impact of two variables: the number of selected frames and the number spatial visual tokens per frame. More specifically, we investigate how different combinations of these variable affect the performance of a trained model. Interestingly, as shown in Figure 3(a) and  3(b), we observed that, compared to the maximum setting with 16 selected frames and 32 spatial visual tokens per frame, it exhibits only a marginal performance drop when the number of selected frames is reduced while maintaining a sufficient number of visual tokens. For example, when the number of frames is set to 16, reducing the number of visual tokens from 16 to 8 results in a performance drop of 0.6% and 1.4% on MVBench and VideoMME, respectively. In a more extreme setting, reducing the number of visual tokens from 8 to 4 results in a performance drop of 3.5% and 2.2% performance drop on MVBench and VideoMME, respectively. We argue that this is because 8 key frames are sufficient for most video understanding tasks. When adjusting the number of spatial visual tokens, we observe only trivial performance drops when reducing the number from 32 to 16 and from 16 to 8, with drops of 0.8% and 0.6% on MVBench and 1.0% and 0.9 % on VideoMME, respectively. However, when the number spatial visual tokens is further reduced, such as from 8 to 4, a significant performance drop of 3.5% and 2.2% on MVBench and VideoMME, respectively. We argue that insufficient spatial visual tokens essentially lead to a lack of spatial perception, resulting in a significant performance drop. Furthermore, the aforementioned observations suggest that setting the number of spatial visual tokens and frames to 8 each achieves an optimal balance between performance and computational efficiency in our method, which still significantly surpasses LLaMA-VID.
Lastly, since the number of selected frames does not affect performance on image benchmarks, we report only the performance changes associated with different numbers of visual tokens. As depicted in Figure 3(c), we observe a trend where an increase in the number of visual tokens leads to better performance on image benchmarks; however, the performance gain becomes marginal as the number of visual tokens approaches 16. The trend observed in image benchmarks suggests that the potential of spatial visual tokens in B-VLLM remains underutilized and a potential improvement direction for B-VLLM.
5.4 Visualization
@@
As shown in Figure 4, we visualize the frame selection and spatial visual tokens for two short-medium videos and one long video. In the first case, the frame selection module locates the frame that contains the banana taped on the wall. Then based on the question, the visual tokens near the banana and tape are sampled and used by LLM, Lastly, the LLM backbone answers the question correctly based on the selected visual tokens. In the second video, the frame selection module samples frames from both ends of the video, showing our method selects frames based on the question rather than through random or uniform sampling. In contrast, for the case of a long video where the question pertains to the entire video, the selected frames are distributed across the video rather than clustered at specific points.
6 Limitation
We acknowledge that our proposed method has limitations in handling multi-round conversations based on the same video. In the video-based multi-round conversation, B-VLLM repeats the frame selection and spatial visual token procedure as the context shifts during conversation, resulting in additional computation cost. Moreover, as discussed in Section 5.3, the spatial visual tokens in B-VLLM are underutilized, and the full potential of our method for image data has yet to be fully explored. Additionally, the Gumbel-Softmax sampling technique may disrupt the temporal order of selected frames, potentially compromising our method’s performance in temporal perception. Addressing these limitations could further enhance the capabilities of VLLMs.
@@ 7 Conclusion
We present a novel vision LLM framework - B-VLLM, designed to address the issue of visual token overload in video understanding with LLMs. Our approach controls the number of visual tokens by adaptively selecting the most relevant frames and applying dynamic spatial token sampling and merging. Comprehensive experiments show B-VLLM achieves SOTA performance on various video benchmarks.

### v2

@@ 4 Implementation
[…]
B-VLLM 32 71.9 84.5 45.1 1455 76.6 74.5 58.4 54.8 67.1
5 Experimental Result & Discussion
@@ 5.1 Video Benchmark Quantitative Evaluation
We evaluate the proposed B-VLLM on various video benchmarks in a zero-shot manner. The video duration in the these benchmarks spans from 10 second to 1 hour, where the details are summarized in the supplementary materials. For open-end answering, we have MSVD-QA [55], MSRVTT-QA [55] and ActivityNet-QA [12], which aim to evaluate the capacity of video captioning. For multi-choice question answering (MCQA), we have MVBench [27] and VideoMME [15], which are multi-purpose video understanding benchmarks; EgoSchema [40] focuses on evaluating the temporal understanding capability; Perception [43] aims to examine reasoning and perception skills of VLLMs; and VNBench [60] evaluates a VLLM using the needle in a haystack manner: asking questions about one or more irrelevant images that are randomly inserted into a video.
@@
Quantitative results for open-end question answering and MCQA benchmarks are reported in Table 1. Overall, B-VLLM achieves SOTA performance on MSVD-QA, Perception, VNBench, Egoschema, and VideoMME. Specifically, when integrating with existing VLLMs, it significantly boosts performance to the two baselines across most video benchmarks. Notably, our method achieves substantial improvements in long-video scenarios. For example, VideoMME-Long evaluates VLLMs performance in extremely long-video settings, with an average video duration of 2466.7 seconds. Our proposed method introduced an 10.5% performance gain for LLaMA-VID [30] and 3.8% for VideoLLaMA2 [9]. Context window length is one of the key factors limiting existing VLMs to handle long videos. For example, LLaMA-VID encodes each frame with 2 tokens, it struggles with extremely long video due to its context window limit and potential key information loss. Our adaptive frame selection module addresses this limitation by selecting a fixed number of frames most relevant to the given context, ensuring the number of visual tokens remains within the context window of LLaMA-VID. As a result, LLaMA-VID consistently performs well in medium and long video settings. We also observe a noticeable performance drop when VideoLLaMA2 handles medium and long videos, which we attribute to its uniform frame sampling strategy that may overlook the frames relevant to the context. The integration of our method mitigates this issue by filtering out irrelevant frames, leading to 1.7% and 3.8% performance gains for VideoLLaMA2.
@@
Additionally, the proposed method remains effective in the short-video settings. For example, LLaMA-VID achieves a 4.5% and 5.8% performance gain in MVBench and Perception through the integration of our method. Similarly, in open-end benchmarks, performance improvements of 1.7%, 0.2% and 1.5% are observed in MSVD-QA, MSRVTT-QA and ActivityNet-QA, respectively. This substantial improvement can be attributed to the enhanced spatial information from the increased number of spatial visual tokens compared to LLaMA-VID. Similar performance improvements are also observed in Video-LLaMA2, with gains of 1.0%, 2.4%, 1.8%, 2.1%, 1.9%, 3.3% and 0.9% on MVBench, Perception, VNBench, Egoschema, MSVD-QA, MSRVTT-QA and ActivtiyNet-QA, respectively.
@@
An interesting observation is that VLLMs employing uniform sampling achieve better performance in short-video scenarios, particularly when the number of sampled frames closely matches the video length in seconds. For example, VideoChat2 [27] ranks highest on MVBench (16 frames), and LLaVA-Next-Video [59] ranks second on Perception (23 frames). Although these methods could theoretically improve performance in longer videos by sampling more frames, they quickly exceed the context-window limitations, increasing computational costs and reducing performance. In contrast, our method maintains consistent performance across short and long videos due to effective frame selection and spatio-temporal merging techniques. Refer to the supplementary material for more experimental result.
@@ 5.2 Image Benchmark Quantitative Evaluation
Our proposed method yields comparable, if not superior, performance, despite not being specifically designed for the image-only setting and using significantly fewer visual tokens per image than existing image-based VLLMs. In particular, B-VLLM outperforms classic image-only VLLM, LLaVA-1.5 [33] on MMBench [34], ScienceQA-Image [36], and SEED-Image [25] by 7.1%, 7.7%, and 1.0%, respectively. Notably, comparable results are observed on POPE [29] and MME [13], despite LLaVA-1.5 [33] consuming 18 times more visual tokens than ours. These benchmarks primarily focus on reasoning and general image understanding, where fine-grained visual details are less critical and typically require fewer visual tokens. However, as noted by LLaMA-VID [30], more visual tokens generally lead to better performance on certain image-only benchmarks. The limitation in the number of visual tokens compromises our method’s performance on VizWiz [17], VQAv2 [16] and TextVQA [49], where these benchmarks examine spatial reasoning and OCR. For example, TextVQA [49] focuses on detecting text within image, while GQA [20] emphasizes complex scene understanding.
In terms of performance gain, integrating our proposed method results in significant improvements for LLaMA-VID [30] and VideoLLaMA2 [9] on image benchmarks. For LLaMA-VID, we observe gains of 9.5% and 10.2% on MMB and VQAV2; and for VideoLLaMA2, non-trivial gains can be observed when incorporating with our method. We believe these gains stem from the effectiveness of the visual token sampling module, as it dynamically selects the most informative tokens, reducing the overall token count while maintaining performance.
 F.S. T.M. $S.$ MVB VMME MMB POPE
@@
Table 3: Ablation study on adaptive frame selection (F.S), temporal frame token merging (T.M.), and spatial visual token sampling (S.S). Experiments are done on LLaMA-VID.
(a) Impacts on the number of frames selected LL* and spatial vis. tokens $R$
(b) Impacts on the temperature $\tau$ and frame similarity threshold $\gamma$
(c) Impacts on maximum number of token $\theta$
Figure 3: Impacts on the hyperparameters.
@@ 5.3 Ablation Study
Effectiveness of the proposed modules in terms of their performance is reported in Table 3. In this experiment, LLaMA-VID [30] is selected as the base model. MVBench [27] and VideoMME [15] are used to evaluate video understanding performance, while MMBench [34] and POPE [29] are used to assess image understanding performance. Starting with the baseline, without the integration of our proposed modules, LLaMA-VID [30] displays moderate performance on the short-video benchmark MVBench [27]. However, compromised performance can be observed on the medium-to-long video benchmark Video-MME [15] due to the limitation of the context window length. In addition, LLaMA-VID [30] typically performs worst on image benchmarks, due to its poor spatial perception caused by a deficit of spatial visual tokens. With the introduction of the text-conditioned adaptive frame selection (F.S.) module, the baseline model no longer suffers from its own token selection mechanism, resulting in 2.9% performance boost on the medium-to-long video benchmark VideoMME [15]. With the temporal frame token merging (T.M.) mechanism, the model mitigates temporal redundancy in duplicated frame selection, resulting in noticeable performance gains on both short and medium-to-long video benchmarks. Similarly, the integration of spatial visual token sampling (S.S.) enables dynamic adjustment on the number of spatial visual tokens to balance the temporal and spatial cues, resulting in performance gains 1.4% and 1.5% on MVBench and VideoMME, respectively. More ablation experiments can be found in supplementary material.
Ach./Token. Time MVB EgoS Perp. VMME
Random - 27.7 20.0 33.0 25.0
Q-Former 25 hrs. 46.5 44.3 48.0 44.5
Resampler 23 hrs. 44.3 45.5 46.8 44.0
Mean Pooling 19 hrs. 41.2 40.0 41.1 37.6
CLS Token 10 hrs. 34.5 34.1 37.8 36.8
Table 4: Q-Former v.s Resampler in Frame Selection and Mean Pooling v.s [CLS] tokens in VLLMs.
@@
Impacts on the number of frame selected LL* , spatial visual tokens $R$ and maximum visual token count $\theta$ are illustrated in Figure 3(a) and  3(c). B-VLLM is selected as base model in the upcoming experiments.First, we examine the impact of two variables: the number of selected frames LL* and the number spatial visual tokens $R$ per frame. More specifically, we investigate how different combinations of these variables affect the performance of a trained model. Interestingly, as shown in Figure 3(a), we observed that, compared to the maximum setting with 32 selected frames and 32 spatial visual tokens per frame, it exhibits only a marginal performance drop when the number of selected frames is reduced while maintaining a sufficient number of visual tokens. For example, when the number of spatial visual token $R$ is set to 32, reducing the number of selected frames LL* from 32 to 16 results in a performance drop of 0.9% and 1.3% on MVBench and VideoMME-Medium, respectively. In a more extreme setting, reducing the number of frame selected LL* from 16 to 8 results in a performance drop of 0.5% and 0.4% performance drop on MVBench and VideoMME-Medium, respectively. We argue that this is because 8 to 32 key frames are sufficient for most video understanding tasks. When using 32 frames, dropping the number of visual tokens $R$ from 32 to 16 results in performance drops of 0.6% on MVBench and 0.4% on VideoMME-Medium. This indicates that current video understanding tasks tend to rely more on temporal dynamics rather than on detailed spatial information. Secondly, we examine how the maximum token count $\theta$ affects the trained model while keeping both the number of selected frames LL* and the number of spatial visual tokens $R$ fixed at 32. As shown in Figure 3(c), increasing the token count generally improves performance, but the gains for B-VLLM plateau at 512 tokens. This indicates B-VLLM can deliver strong performance without a high computational budget.
Impacts on temperature $\tau$, similarity threshold $\gamma$ are presented in Figure 3(b). In these experiments, the number of selected frames LL* and spatial visual tokens $R$ are both set to 32, and the maximum token count $\theta$ is fixed at 768. Our results indicate that the optimal performance is achieved with a lower $\tau$ value in combination with a higher $\gamma$ value. Where $\tau$ governs the discreteness of Gumbel-Softmax during frame selection and higher $\gamma$ values prevents dissimilar frames from being merged. For example, when $\tau$ is set to 0.1, reducing $\gamma$ from 0.75 to 0.5 results in 1.2% and 1.7% performance drop on MVBench and VideoMME-Medium, respectively. Similarly, when $\gamma$ is fixed to 1.0, lowering $\tau$ from 0.1 to 0.5 results in 0.7% and 1.2% performance drops. We find that a high $\gamma$ prevents merging dissimilar frames, while a low $\tau$ mitigates the oversmoothing of frame features caused by the Gumbel-Softmax.
Q-Former v.s. Resampler in Frame Selection. Besides Q-Former, we also verified with other commonly used frame selection module - Resampler, where the results are reported in Table 4. Resampler is originally designed for visual token abstraction, for fair comparison, we provide it with both visual and text tokens, setting the same number of query tokens as in Q-Former. Notably, Q-Former outperforms Resampler on most selected benchmarks. We argue that the Resampler’s reliance on self-attention mechanism limits its performance in processing cross-modality tokens.
Efficacy of [CLS] token. To validate the efficacy of [CLS] tokens in frame selection, we compare its performance against using visual tokens with mean pooling as frame features. For simplicity, we conduct preliminary experiment on VideoLLaMA2 [9] where the modality projector is replaced with a single MLP layer. It is worth mentioning that, none of the proposed module is present in this preliminary experiment. All experiments are trained using the same settings, with 8 frames selected from each video. As shown in Table 4, [CLS] tokens enables the model to consistently outperform random guessing while being more training-efficient than mean pooling. This efficiency is preferable and the effectiveness is sufficient for frame selection, as the goal is to filter out most irrelevant frames in a coarse stage. Refer to the supplementary material for more details.
5.4 Discussion
@@
Visualization. As shown in Figure 4, we visualize frame selection and spatial tokens for two short-medium videos and one long video. In the first case, the frame selection module locates the frames with the banana taped to the wall. Then based on the question, the visual tokens near the banana and tape are sampled and used by LLM, Lastly, the LLM backbone answers the question correctly based on the selected visual tokens. In the second video, the selected frames are from both ends, demonstrating question-based selection rather than random or uniform sampling. In contrast, for long videos with overarching questions, the frames selected are evenly distributed. Limitation. We recognize that using [CLS] tokens for frame selection has its limitations since pooled visual tokens are more informative, as an alternative strategy for exploring more accurate frame selection. See the supplementary material for further discussion.
@@ 6 Conclusion
We present a novel vision LLM framework - B-VLLM, designed to address the issue of visual token overload in video understanding with LLMs. Our approach controls the number of visual tokens by adaptively selecting the most relevant frames and applying dynamic spatial token sampling and merging. Comprehensive experiments confirms the effectiveness of B-VLLM. We also highlight the limitations of our method to inform future directions.
Acknowledgements
This work was supported by the Australian Research Council (ARC) Linkage Project #LP230100294, ARC Discovery Project #DP210102674 and Edith Cowan University Science Early Career and New Staff Grant Scheme.
Supplementary Material
Table 5: Statistics of training datasets.
 Stage Modality Sample Source
LLaMA-VID [30]
PT Image 558K  [47, 33]
Video 232K [5, 30]
  FT Image 625K  [16, 20, 23, 24, 41, 42, 45, 48]
Video 98K  [12]
Text 40K  [1]
VideoLLaMA2 [9] - Valley [38]
PT Image 558K  [47, 33]
Video 702K [38]
  FT Image 665K  [33]
Video 100k  [39, 38]
Table 6: Statistics of benchmark datasets. Type indicates the QA question type: Open represents open-ended answering and MCQA denotes multi-choice question answering.
 Benchmark Avg. Len (s) # QA Pair Type
MSVD-QA 9.8 10K Open
MSRVTT-QA 15.2 50.5K Open
MVBench 16 4K MCQA
Perception-Test 23 2.7K MCQA
Next-QA 44 4000 MCQA
VNBench 54 5.6K MCQA
VideoMME-S 80.7 0.9K MCQA
ActivityNet-QA 111 8K Open
EgoSchema 180 5K MCQA
VideoMME-M 515.9 0.9K MCQA
VideoMME-L 2466.7 2.7K MCQA
7 Additional Implementation Details
7.1 Training Details
We adopt a two-stage training strategy [33, 30, 9], dividing training into pretraining for modality alignment and fine-tuning for instruction tuning. During pretraining, the learning rate is set to 1e-3 with a linear learning rate schedule, and the total batch size is set to 256. Only the frame selection module and the MLP projection layer, which maps the visual token features to the LLM features, are updated. In the fine-tuning stage, the entire framework is unfrozen except for the preliminary visual encoder. Additionally, LoRA [18] is enabled with the rank set to 128 and $\alpha$ set to 256. The total batch size was set to 128, and the learning rate is 1e-4 with cosine scheduling. Supervision is provided by minimizing cross-entropy for masked text tokens, and the optimizer for both stages is AdamW [35], with DeepSpeed ZeRO2 [3] enabled for memory efficiency. Training is conducted with 16 Nvidia L40S GPUs. Following previous works [27, 39, 9], videos are downsampled to 1 fps in both training and evaluation.
7.2 Dataset for Training
The following provides details on the training datasets utilized in LLaMA-VID-Dataset [30] and VideoLLaMA2 [9]. The dataset used by LLaMA-VID [30] consists of 790K samples for pre-training stage and 763K samples for fine-tuning stage. Compared to the dataset used by LLaMA-VID, Valley [38] contains significantly more video samples for pre-training. There are 1.25M samples for pre-training stage and 765K sample for fine-tuning stage. The difference in the volume of video samples may explain the performance differences observed between LLaMA-VID and VideoLLaMA2.
7.3 Video Benchmark Summary
The details of the video benchmarks are summarised in Table 6. We selected benchmarks with diverse durations, avoiding a focus on a fixed average video length. For example, the MSVD-QA [55] benchmark consists of short videos with an average duration of 9.8 seconds. In contrast, the VideoMME-L [15] is a representative of long video benchmark, with an average duration of 2466.7 seconds and a maximum video length capped at 1 hour. We argue that our selection of video benchmarks reflects the diversity of video durations in real-world scenarios, demonstrating the applicability and robustness of our model in practical environments.
7.4 B-VLLM Experimental Setup
Our B-VLLM is implemented based on VideoLLaMA2 [9], where Qwen2 [56] is selected as the backbone LLM. Each frame is encoded as 768 preliminary visual tokens using EVA-CLIP [sun2023eva], Subsequently, 32 frames are selected using the frame selection module and temporal frame token merging module. The visual tokens corresponding to the selected frames are encoded into 32 spatial visual tokens by our spatial visual token sampling module.
8 Additional Experiment & Discussion
8.1 Additional Discussion on Different Frame Selection Features.
As reported in Table 7, we additionally report the performance of selecting frames by using three other feature extraction methods: (a) max pooling (b) mean pooling and (c) Qformer. To investigate the impact of these variants in B-VLLM’s fine-grained spatial perception, we also report the performance of OCR and Counting in VMME, which are more relevant to spatial capability. Though [CLS] token may not be the optimal choice in terms of performance, as discussed in Section 3.2, the rationale of using [CLS] token lies in balancing computational cost and effectiveness.
Table 7: Additional ablation study on frame selection features.
Model Train. MVB EgoS VMME OCR Counting
Mean 13.1h 51.3 51.9 54.4 48.2 36.6
Max 13.1h 49.5 51.6 50.7 45.3 38.1
Qformer 14.0h 52.4 52.9 51.6 41.7 35.8
CLS 10.9h 50.8 51.9 52.9 46.0 34.0
8.2 The Impact of LL*.
As shown in Fig 5, we report additional experiments to investigate the impact of LL* on Short ($\leq$ 3min), Medium ($\leq$ 15min) and Long video ($\leq$ 60 min) settings. As shown below, B-VLLM’s performance generally saturated at L=28L*=28, though it was trained with L=32L*=32. This indicates that there could be an optimal setting for LL*. We argue that this is due to most tasks typically rely on a small portion of frame regardless of the duration of the video. However, we acknowledge that increasing LL* is potentially beneficial for more extreme setting and this review inspired us to develop a LL*-free design as one of our future work.
Figure 5: Additional experiment on the impact of LL*.
8.3 Model’s Parameter Size Comparison
In Table 8, we report our B-VLLM’s parameter size below along with TFLOPs and GAMCs, which are similar to most existing methods.
Table 8: Model size compared with existing methods
Model TFLOPs GMACs Param.
InternVL2 1.83 910 7+1.1B
InternVideo2 1.83 915 7+1.5B
QwenVL 1.92 915 7.7+1.9B
Qwen2VL 1.82 905 7.6+0.6B
VideoLLaMA2 1.82 910 7.6+0.8B
B-VLLM (Ours) 1.81 908 7.6+1.2B
Table 9: Additional Comparison with Recent SOTA methods
Model #Train.Data MVB EgoS Perp. VMME
InternVideo2 39M 60.3 55.8 53.0 41.9
InternVL2 20M+ 66.4 - - 54.0
LLaVA-OV 9.3M 56.7 60.1 57.1 58.2
Qwen2VL - 67.0 66.7 62.3 66.3
B-VLLM (Ours) 2M 50.8 51.9 52.1 52.9
8.4 Additional Comparisons with Recent SoTA.
We provide additional comparisons with several advanced VLLMs in Table 9. Note that these advanced VLLMs are generally trained with at least 5 times more data than our B-VLLM, and these data are often private. This indicates the great potential of our B-VLLM when it is trained with large scale data.
8.5 Inference Speed & Latency
In Table 10, we compare the inference speed between Qwen2VL and B-VLLM on three types of videos: 4-minute video, 8-minute video, and 12-minute video. As reported below, B-VLLM demonstrates clear advantage on inference speed. Specifically, it takes 5.6s for B-VLLM to generate a response for an 8-minute video while taking 24.6s for Qwen2-VL to respond.
Table 10: Inference Speed and Latency
Model 4 Min 8 Min 12 Min
Qwen2VL 6.94s 24.6s 26.2s
B-VLLM (Ours) 3.16s 5.6s 7.9s
8.6 Additional Discussion on Limitations
Limitation. On top of the discussed limitation in the main manuscript, we acknowledge that our proposed method has limitations in handling multi-round conversations based on the same video. In the video-based multi-round conversation, B-VLLM repeats the frame selection and spatial visual token procedure as the context shifts during conversation, resulting in additional computation cost. Moreover, as discussed in manuscript, the spatial visual tokens in B-VLLM are underutilized, and the full potential of our method for image data has yet to be fully explored. Additionally, the Gumbel-Softmax sampling technique may disrupt the temporal order of selected frames, potentially compromising our method’s performance in temporal perception. Addressing these limitations could further enhance the capabilities of VLLMs. We recognize that using [CLS] tokens for frame selection has its limitations since pooled visual tokens are more informative, as an alternative strategy for exploring more accurate frame selection. More limitations are discussed in supplementary material. Additionally, the fixed value $L^{*}$ can limit B-VLLM’s performance on extreme long videos, especially when the number of relevant frames surpasses $L^{*}$.
9 More B-VLLM Qualitative Examples
This section provides additional conversation examples with B-VLLM based on videos. Note that the presented videos are accessible only through the specified source. The examples are depicted in Figures 6 to 10. As shown in Figure 6, when a game trailer is fed into B-VLLM, it accurately identifies the game’s title displayed at the beginning of the video and effectively summarizes the gameplay content, demonstrating B-VLLM’s comprehensive understanding of both temporal and spatial information. In the example shown in Figure 9, during the conversation, B-VLLM accurately identifies the man’s actions. Even though the basketball appears only briefly at the beginning of the video, B-VLLM successfully captures and localizes the nuanced spatial details. As depicted in Figure 10, beyond accurately describing the content of the given video, B-VLLM demonstrates the ability to estimate the age of the man featured in the video, highlighting the spatial perception capabilities of our method.
Figure 6: Visualization on understanding a 2-minutes game trailer containing intensive fighting scenes.
Figure 7: Visualization on understanding a 1-minute AI generated video of a woman in Japan.
Figure 8: Visualization on perceiving a 4-minute wonderful music video containing scenes of singing and dancing.
Figure 9: Visualization on perceiving a 1-minute dancing video a man showcasing his outstanding dancing and basketball skills.
Figure 10: Visualization of understanding a 3-minute video featuring a Kung Fu master demonstrating his martial arts expertise.